A 100kg elevator carrying 4 people with average massses of 55 kg is accelerating upward at 0.8 m/s ^2. What is the tension in the cable?
Answer #1

since the elevator is going up, the pseudo force will act downwards = ma = 100*0.8 = 80 N

T = mg + ma

T = 55*4*9.8 + 80

= 2236 N

so the tension in the cable is 2236 N